Latest update: Tue Feb 2 16:15:10 1999 EST

ACIS Home Page Status Plan Events Test Notes Morgue Anomalies Phone Numbers
Live Engineering Data Redondo Reporter Procedures
Long Form Functional Short Form functional

Current Status:

  • Observatory is in the Shipping Container

    Short Term Plan:

    Test Event History

    Day Activity Results Comments
    1/30/98 Arrive at TRW Did that
    2/2/98 Incoming Short Form Functional Went fine IP&CL 5.1 OK, too
    2/5/98 Mate ISIM to Telescope
    Scheduled 2/6
    2/16/98 Long Form Functional ACIS is fine System has some problems
    2/27/98 Pump out Detector Housing Scott Anderson Productions About 4 torr in DH
    3/7/98 Acoustic Test
    Scheduled 3/5
    3/24/98 Long Form Functional ACIS is fine System has some problems
    3/25/98 Pump out Detector Housing Scott Anderson Productions About 4 torr in DH.
    4/18/98 EMC Test No effects seen at ACIS Scheduled 4/18
    5/16/98 T/V Test Begin Pumping started 0630 Scheduled 5/10
    5/16/98 2.2: Initial pwr. on/bake started 1800
    Internal pressure was 5.5 torr.
    4hr predict -> 11hr actual
    5/17/98 3.1: Transition to Hot #1 started 0500 20hr predict
    5/18/98 3.2: Transition to Cold #1 started 0100 48hr predict
    5/20/98 3.3: Transition to Hot #2 started 0100 32hr predict
    5/21/98 3.4 Transition to Cold #2 started 1940 48hr predict
    5/23/98 3.5: Transition to Hot #3 started 2330 32hr. predict
    5/25/98 3.6: Transition to Cold #3 started 0800 48hr. predict
    5/27/98 4.1: Transition to Hot #4 started 0600 32hr. predict
    5/28/98 4.2: Hot Functionals started 1500
    ACIS shrouds filled with LN2.
    a six day event
    6/5/98 4.3: Hot Thermal Balance
    a two day event
    6/7/98 4.4: Transition to Cold/Eclipse started 2100
    60hr predict
    6/10/98 4.5: Cold Functional started 1900

    6/16/98 4.6: Cold Thermal Balance cut short after 18 hours
    6/18/98 4.7: Stray Light
    Door problems killed the test.
    6/18/98 5.0: Transition to Hot Bake/close valves & door
    6/21/98 Break Vacuum

    6/27/98 Disintegrate ISIM ACIS off and heater inspection
    8/3/98 Reintegrate ISIM
    Short Form OK
    8/10/98 Practice ISIM TV Always in bakeout mode 4 days under hard vacuum.
    8/14/98 Pump Down Started at 2300 PDT
    8/15/98 Start Cold Soak Started at 1600 PDT
    8/24/98 OBF Test Ran fine
    8/24/98 Long Form Partial; Ran fine
    8/29/98 3 Door Cycles
    Worked fine
    8/29/98 Stray Light
    No stray cats, either
    9/2/98 Repress

    9/10/98 CAT-2 Long Form Ran OK Heater Bus Anomaly
    9/11/98 BER Test 20 Cycles
    Mission Sequence Test

    OCC Door Opening Drill

    Extended Systems Test
    Not a problem for ACIS
    Extended Systems Test II

    Test Notes from Engineering

    • ACIS is happy, give or take our (continuing) lack of knowledge as to why the door failed to open at the Observatory-level TV.
    • The fuse/solder joint failure on the PSMC filter board is stable and open.

      Anomalies, Major and Minor:

      • As we come in the door at TRW, the following discrepencies exist:
      • Upon S/C power up, the PSMC sets multiple (command) bits. This is related to a slow, manual ramp up of the voltage. It is definitely dangerous, though we have so far only seen heater functions enabled. The spacecraft folks looked at their fuzes and concluded that they could, after all, stand the surge of a normal, fast power on.
      • It is well known (!) that the 32Kbps hardline command link is noisy, and therefore cannot be used for large sequences of commands without error. During the post-acoustic Long Form it was further discovered that the 2Kbps RF link can't handle command packets larger than approximately 46 16-bit words. Later tests showed that the problem is a lack of buffering in the TCTS command feed. Modifiying our software to rate limit commands below 2Kbps solves the problem.
      • The command and telemetry environment at the Observatory level is no where near as error-free as one might hope. To what extent the problems reside on the ground, and to what extent they are endemic to the flight hardware is impossible for the external observer to ascertain.
      • The door failed to open for the Stray Light Test. All indications are that the stuck mechanism (actual failure mechanism unknown) resulted in bursting the rupture disk of the "open" actuator. A summary report is available. as well as a discussion about how one might do commanding differently.
      • Applying the new and improved door opening procedure op cit we managed to confuse ourselves mightily. A discussion about how not to operate a Starsys actuator on ACIS is provided for the amusement of the reader. After trying both the pulse mode and the single-135-second mode, one can conclude that no current proposal is satisfactory for on-orbit use.
      • After a completely successful Long Form Functional as part of the (ambient) Observatory level CAT-2 test, we cycled the Detector Assembly heaters on. Thus did we observe that the input voltage level within the PSMC was not perfectly healthy.
      • During the Enhanced Systems Test it was discovered that the AXAF RF receivers don't like to see more than 29 binary zeros in a row -- which happens is almost every parameter block we send up, for instance. The receiver drops off line and the command gets lost.

        Phone Numbers

        TRW ACIS R4T office is: 310-813-5816 (has voice mail)
        TRW AXAF R4T FAX stuff: 310-812-2788
        TRW ACIS R7A EGSE area: 310-812-2200 or 2-5070
        TRW ACIS M4 EGSE area: 310-812-1174
        TRW HRC M4 EGSE area: 310-813-7301
        TRW STC M4 EGSE area: 310-813-4514
        Residence Inn: 310-546-7627

        Name Function Work Home Notes
        Demitrios Athens Software GSE 258-5285 617-566-7668
        Mark Bautz Proj. Science 253-0023 781-863-2024
        Ed Boughan System Eng. 253-6146 508-470-1043 pager: 800-439-9827
        Mike Doucette Electrical GSE 253-3571 617-438-7690
        Peter Ford Software 253-7277 617-661-0868
        Bob Goeke Project Eng. 253-1910 781-729-8484 pager: 800-453-3658
        Fred Kasparian Mechanical 253-1065 617-327-6330
        Brian Klatt Quality 253-7555 508-785-1511
        Bill Mayer Project Mgr. 253-7552 508-692-3318
        Jim O'Connor Manufacturing 253-6139 617-335-2915
        Ellen Sen Thermal 253-8342 781-665-3568
        Matt Smith Contamination 252-1736 508-435-9970
        Mark Bautz Project Science 253-0023 617-566-6320
        Fred Baganoff pblock Guru 253-6892 781-643-4648
        Beverly LaMarr Science Coordinator 258-8153 617-776-0621 Lab: 3-7466
        Scott Texter TRW/Integration 310-813-1904
        pager: 888-975-7157
        Eric Vrem TRW/SIM 310-813-9075

        Joe Vogrin TRW/Thermal 310-813-0478

        Phase I T/V test record is still available.

        Phase II T/V test record is still available.

        ISIM Integration record is still available.

        Page maintainer: Bob Goeke

        There have been 168 accesses to this page since 9/14/2022.